OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

look at the last 2 terms, 72 and 24. Is there anything you can divide 72 by to get 24? If there is, is it the same thing you divide 216 by to get 72? If so, you've found the rule for this sequence and can find the next 3 terms below 24.

OpenStudy (anonymous):

@JakeV8 would it be 12 for 72 and 24?

OpenStudy (anonymous):

@JakeV8 72 for 216 and 72? or smaller

OpenStudy (anonymous):

or divide all by 12?

OpenStudy (anonymous):

72 / 12 = 6, so no... try 72 / 3 72 / 3 = 24 Now check: 216 / 3 = 72 The way I realized it was dividing by 3 was to switch it around and divide 72/24 to see what I got... when I got 3, I tried dividing 216 by 3 to see if I got 72. You are looking for that rule of "what do I divide each term by to get the next term?"

OpenStudy (anonymous):

648 / 3 = 216 216 / 3 = 72 72 / 3 = 24 So, your task is: 24 / 3 = ?? (next term after 24) then use that answer, divide by 3 to get ??? (2nd term after 24) finally, use that ??? and divide by 3 to get the last term (3rd term after 24)

OpenStudy (anonymous):

The rule on this problem is "divide by 3 to get next term" but you need to learn to spot the rule in whatever the sequence is

OpenStudy (anonymous):

24/3 = 8 8/3= 2.66 2.66/3 0.88 ???

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

648 * 1/3 * 1/3.....

OpenStudy (anonymous):

yes :) but I would just leave it as fractions, unless you've been told decimals are required. 8, 8/3 8/9 (because 8/3 divided by 3 is 8/9)